# sPhinX: Sample Efficient Multilingual Instruction Fine-Tuning Through N-shot Guided Prompting

### v2

@@ Abstract
Despite the remarkable success of LLMs in English, there is a significant gap in performance in non-English languages. In order to address this, we introduce a novel recipe for creating a multilingual synthetic instruction tuning dataset, sPhinX, which is created by selectively translating instruction response pairs from English into 50 languages. We test the effectiveness of sPhinX by using it to fine-tune two state-of-the-art models, Phi-3-small and Mistral-7B and then evaluating them across a comprehensive suite of multilingual benchmarks that test reasoning, question answering, and reading comprehension. Our results show that Phi-3-small and Mistral-7B fine-tuned with sPhinX perform better on an average by 4.2%pt and 5%pt respectively as compared to the baselines. We also devise a strategy to incorporate N-shot examples in each fine-tuning sample which further boosts the performance of these models by 3%pt and 10%pt respectively. sPhinX outperforms other multilingual instruction tuning datasets on the same benchmarks along with being sample efficient and diverse, thereby reducing dataset creation costs. Additionally, instruction tuning with sPhinX does not lead to regression on most standard LLM benchmarks.
@@ 1 Introduction
Large Language Models (LLMs) have been shown to perform several tasks exceptionally well in English, however, performance in some non-English languages still lags behind Ahuja et al. (2023); Asai et al. (2023). Further, the gap between the performance of Large Language Models (LLMs) and Small Language Models (SLMs) is more pronounced Ahuja et al. (2024) on
@@
non-English languages. Cui et al. (2023) and Balachandran (2023) follow the approach of training the models further on datasets for specific languages, however this can lead to catastrophic forgetting and hurt the performance on English Zhao et al. (2024). Several techniques have been proposed to bridge this gap, such as incorporating better pre-training data in various languages and enhancing base tokenizers Xu et al. (2024); Dagan et al. (2024). However, most of these changes need to be implemented at the pre-training stage which demands extensive data and computational resources, making it practically unfeasible in many scenarios Brown et al. (2020). Consequently, the most well-studied technique involves fine-tuning models for specific languages and tasks. Instruction fine-tuning (IFT) has become a popular technique for enhancing the performance of language models in specific languages. This method combines the benefits of both the pretrain–fine-tune and prompting paradigms Wei et al. (2021).
@@
Sample diversity is crucial for instruction tuning in multilingual datasets. Many recent datasets are created by translating English content into other languages or using self-instruct techniques from seed prompts Li et al. (2023); Taori et al. (2023), both of which can limit diversity. Machine translation can lead to loss of semantic information Baroni and Bernardini (2006), while self-instruct methods may produce repetitive and homogeneous samples Wang et al. (2022). This underscores the importance of having a dataset with diverse set of samples.
Aggarwal et al. (2024) evaluate several models fine-tuned using Parameter Efficient fine-tuning and find that there is a gain in multilingual performance across several SLMs for many low-resource languages, with some high-resource languages performing worse after fine-tuning. However, the performance gains often do not match the performance of larger models, such as GPT-4 and Gemini and can be inconsistent across languages. Hence, there is a need to study instruction tuning for better multilingual performance in SLMs.
@@
In this paper, we present a novel recipe for creating a multilingual synthetic instruction tuning dataset, sPhinX.  It comprises 1.8M instruction-response pairs in 51 languages, derived by selectively translating the Orca instruction tuning dataset Mukherjee et al. (2023) using GPT-4 Achiam et al. (2023). We assess the effectiveness of sPhinX by fine-tuning two models — Phi-3-small and Mistral-7B — across a range of evaluation benchmarks that test various language model capabilities like reasoning, question answering, and reading comprehension. We compare models fine-tuned on sPhinX with those using other synthetic multilingual instruction tuning datasets like Aya Üstün et al. (2024), Multilingual Alpaca Taori et al. (2023), and Bactrian Li et al. (2023) and observe significant performance gains.
@@
  • We introduce a novel approach to generate synthetic data for multilingual instruction tuning by selectively translating the Orca dataset with the assistance of GPT-4 (§3.1)
  • We devise LAnguage-Specific N-shot Guided Instruction fine-tuning (LANG) strategy for enhancing the multilingual capabilities of LLMs (§4.2)
  • We conduct an extensive number of experiments on different multilingual instruction tuning datasets to test generalizability in multilingual settings (§5)
  • We also conduct an in-depth analysis where we find sPhinX to be more sample efficient and diverse across languages (§3.3)
@@ 2 Related Work
2.1 Multilingual LLMs
Recently, there has been there has been interest in developing and improving SLMs such as LLaMA Touvron et al. (2023), Mistral Jiang et al. (2023), Phi3 Abdin et al. (2024), and Gemma Team et al. (2024). In a relatively brief period, researchers have developed numerous monolingual and multilingual versions of these foundational models, either by pre-training from scratch on specific languages or groups of languages, or by further fine-tuning on them Nguyen et al. (2023); Gala et al. (2024); Balachandran (2023); Cui et al. (2023); Zhang et al. (2023); Qin et al. (2024). Prior work has shown that smaller models show promising results in high-resource languages but perform worse on low-resource ones Ahuja et al. (2024) compared to larger models. To address this, researchers have proposed using multilingual corpora for pre-training, expanding the vocabulary Wang et al. (2019) and continual training Liu et al. (2023), though this increases training costs significantly. fine-tuning for specific tasks, such as translation, has been explored but often struggles with generalization  Mishra et al. (2021). In contrast, instruction tuning has shown to improve generalization to unseen tasks, enhancing the models’ ability to understand and respond to human instructions  Mishra et al. (2021); Ouyang et al. (2022). This makes the models more versatile and capable of handling a wide variety of tasks.
@@ 2.2 Multilingual Instruction fine-tuning
Early studies focused on fine-tuning pre-trained models on a variety of languages through data augmentation for a single task Hu et al. (2020); Longpre et al. (2021); Asai et al. (2022). Currently, the approach has shifted on fine-tuning these models using a wide-variety of tasks Longpre et al. (2023); Ouyang et al. (2022). Models such as BLOOMZ Muennighoff et al. (2022) and mT0 Muennighoff et al. (2022) makes significant strides in improving the multilingual performance of decoder-based models Ahuja et al. (2023). There have been multiple multilingual instruction datasets and models proposed such as Bactrian Li et al. (2023), AYA Üstün et al. (2024), polyLM Wei et al. (2023b) after BLOOMZ and mT0. However, these models still do not perform as well as English on other languages, with the gap being particularly large for low-resource languages and languages written in scripts other than the Latin script  Ruder et al. (2021); Ahuja et al. (2023); Asai et al. (2023); Ahuja et al. (2024). In this work, we aim to narrow the performance gap by introducing a strategy for creating datasets for multilingual instruction tuning and recipes for fine-tuning, which we will discuss in the following sections.
Figure 1: The above figure describes translating using a Translation API vs Selective Translation
@@ 2.3 Multilingual Synthetic Data Generation
Most instruction-tuning datasets across multiple languages typically focus on general tasks rather than specific reasoning capabilities. While datasets like Orca Mukherjee et al. (2023) and Orca 2 Mitra et al. (2023) exist in English, they highlight a prevalent issue: current methods often prioritize style imitation over leveraging the reasoning abilities found in large foundation models (LFMs). The Orca dataset addresses this by imitating rich signals from GPT-4, including explanation traces and step-by-step thought processes Wei et al. (2023a), guided by assistance from ChatGPT. In order to create multilingual datasets, researchers commonly use translation APIs or LLMs to translate English-specific datasets into target languages. For example, the Bactrian dataset Li et al. (2023) translates Alpaca and Dolly instructions into 52 languages using the Google Translator API and generates outputs with GPT-3.5 turbo. However, these translated datasets often struggle to encode semantic information effectively Baroni and Bernardini (2006). Our dataset approach aims to tackle these challenges by selectively translating only the essential portions of multilingual inputs. This strategy not only preserves semantic information but also accommodates diverse linguistic contexts, thereby enhancing the overall quality and applicability of instruction-tuning datasets across languages.
@@
In this section, we describe our dataset construction methodology (§3.1), dataset filtering and cleaning pipelines (§3.2), and exploratory analysis (§3.3) for determining diversity of our data in comparison to other instruction tuning datasets.
Figure 2: Some examples of input queries and its counterpart existing in the hindi version of the Multialpaca dataset and if it was generated using the Selective Translation strategy
@@ 3.1 Dataset Construction
Mukherjee et al. (2023) illustrate how smaller models can replicate the reasoning processes of LLMs and learn from detailed signals like as explanation traces, step-by-step thought processes, and other intricate instructions obtained from data annotated by ChatGPT and GPT-4111GPT-4 inference hyper-parameters in Azure OpenAI interface set as: temperature=0.0, top_p=1, frequency_penalty=0, presence_penalty=0, stop=[”¡—im_end—¿”, ”¡—im_start|¿”]. Their dataset consists of 5M ChatGPT and 1M GPT-4 generated instruction-response pairs. Inspired by this work, we utilized 1M GPT-4 generated instruction-response pairs of their dataset and construct our dataset along similar lines by selectively translating them into 50 different languages with the help of GPT-4. We categorize them into three groups: high-resource, mid-resource, and low-resource languages as outlined in Table 11. For high resource languages, we randomly sample 100k instruction-response pairs from the Orca 1M dataset and generate the responses from GPT-4 using Selective Translation as shown in Figure 1. Similarly, we leverage the same strategy for medium and low resource languages by sampling 50k and 25k pairs respectively. Although GPT-4 performs competitively with commercial translation systems (Google Translate & Bing Translate) it still lags behind on medium and low resource languages Jiao et al. (2023); Hendy et al. (2023). Furthermore, as highlighted in Chang et al. (2023), fine-tuning with a large set of samples from medium and low-resource languages might lead to catastrophic forgetting of high-resource languages. Therefore, we deliberately create fewer samples for medium and low-resource languages than for high-resource ones.
@@
A fundamental problem with using an off-the-shelf translation API is the lack of semantic and task awareness, in addition to translationese Baroni and Bernardini (2006)), which can result in poor quality training data. Consider for example the task of Machine Translation as part of the instruction, wherein the language of the source sentence should be retained. However, an off-the-shelf API, without task awareness, would translate it, resulting in an ill defined instruction. To mitigate this issue, we prompt GPT-4 to selectively translate the instructions, so that the tasks are translated into the appropriate language without changing the semantic meaning. Figure 2 illustrates this with concrete examples. The first example demonstrates the aforementioned translation inconsistency issue for an instruction asking for a French equivalent of an English phrase. The second example demonstrates the consequence of direct translations in the MultiAlpaca dataset: wherein the translation of the task input results in the task being ill-defined based on the instructions. As demonstrated, our proposed Selective Translation method is able to keep the semantic information of the task intact while translating the instructions. For the exact prompt, please refer to Figure 6 in the Appendix.
3.2 Dataset Filtering and Cleaning
Input: listOfSentences: list of strings
1 Function dataFilter(listOfSentences):
2       englishWords $\leftarrow$ set of English words from NLTK;
3       foreach sentence in listOfSentences do
4             cleanedSentence $\leftarrow$ replace all punctuations, digits, and single characters with a single space;
5             cleanedSentence $\leftarrow$ replace all sequences of whitespace with a single space;
6             wordCount $\leftarrow$ 0;
7             foreach word in cleanedSentence do
8                   if word.lower() in englishWords then
9                         wordCount $\leftarrow$ wordCount + 1;
10
11
12
13            content $\leftarrow$ wordCount / len(cleanedSentence);
14             if content $>$ 0.90 then
15                   remove sentence from listOfSentences
16
17
Algorithm 1 Data Filtering Algorithm
Figure 3: t-SNE Visualisation of 1000 samples equally distributed in 10 languages
After generating these translations, we analyze how many instruction-assistant pairs were unsatisfactorily translated, such as incomplete or missing translations, by identifying the English content within the instances. We use the NLTK [^2] https://www.nltk.org/ library to identify the fraction of English words in the sample. The NLTK corpus contains approximately 240,000 English words, making it suitable for our use case. We manually analyzed various proportions of English content in samples of each language and found that the samples were satisfactory when the English content was within 90% for all languages. If the English content exceeded 90% in a language that uses a different script, those samples were removed from our dataset as shown in our data filtering algorithm here. Finally, we manually inspected a small sample of the dataset to ensure the quality of the translations, and found them to be of good quality.
@@ 1 Function dataFilter(listOfSentences):
After filtering, the final dataset comprised of 1.8 million samples in 51 languages, divided into three subsets: Train, Test, and Few-shot. We partitioned each language’s dataset to ensure the Test and Few-shot sets contained 2000 and 1000 samples respectively, with the Train set consisting the remaining data. This approach guarantees consistent distributions across languages in the Test and Few-shot sets, ensuring equitable representation regardless of training distribution. The train, test, and few-shot sets are in the ratio of 92 : 5.3 : 2.7.
3.3 Data Analysis
A primary issue that we hypothesize (and find preliminary evidence for from basic analysis of prior Multilingual IFT datasets) is the lack of diversity of samples. In order to validate this issue, we use t-Distributed Stochastic Neighbor Embedding (t-SNE) Cai and Ma (2022) to visualize the structure of both prior multilingual IFT datasets as well as sPhinX. We use the model paraphrase-multilingual-MiniLM-L12-v2 from huggingface’s sentence-transformers Reimers and Gurevych (2020), Reimers and Gurevych (2019) library [^3] https://huggingface.co/sentence-transformers/paraphrase-multilingual-MiniLM-L12-v2 to generate the embeddings. The hyperparameters were set as follows: learning rate at 200, perplexity at 30, and number of iterations (n_iter) at 1000. Concretely, we use 1000 samples equally distributed in 10 languages. Figure 3 demonstrates that sPhinX demonstrates greater diversity compared to prior datasets. Particularly, Bactrian and Multilingual Alpaca appear less diverse, which we hypothesize might be due to their generation method involving translations of the same set of task limited English samples into different languages. By virtue of using a more diverse seed set (Mukherjee et al., 2023), we circumvent this issue by design. Furthermore, one notable differentiating factor of sPhinX is the substantial amount of code-switched data the inherent nature of Selective Translation. This in turn enhances the diversity of samples across multiple languages. Finally, we also observe the Aya dataset to be more diverse than other prior counterparts, though it still shows less diversity across task instructions.
Dataset Average Token
Length/Sample
Aya 2240
Bactrian 2465
Multilingual Alpaca 1620
sPhinX-0s 544
sPhinX 3100
Table 1: Average Token Length in each dataset

### v4

@@ Abstract
Despite the remarkable success of large language models (LLMs) in English, a significant performance gap remains in non-English languages. To address this, we introduce a novel approach for strategically constructing a multilingual synthetic instruction tuning dataset, sPhinX . Unlike prior methods that directly translate fixed instruction-response pairs, sPhinX enhances diversity by selectively augmenting English instruction-response pairs with multilingual translations. Additionally, we propose LANGIT, a novel N-shot guided fine-tuning strategy, which further enhances model performance by incorporating contextually relevant examples in each training sample. Our ablation study shows that our approach enhances the multilingual capabilities of Mistral-7B and Phi-3-small  improving performance by an average of 39.8% and 11.2%, respectively, across multilingual benchmarks in reasoning, question answering, reading comprehension, and machine translation. Moreover, sPhinX maintains strong performance on English LLM benchmarks while exhibiting minimal to no catastrophic forgetting, even when trained on 51 languages.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated exceptional performance across various tasks in English. However, their performance in some non-English languages remains comparatively limited Ahuja et al. (2023); Asai et al. (2024). Further, the gap between the performance of Large Language Models (LLMs) and Small Language Models (SLMs) is more pronounced (Ahuja et al., 2024) in non-English languages. Cui et al. (2023) and Balachandran (2023) utilize the method of fine-tuning models on datasets focused on particular languages. However, this can lead to catastrophic forgetting, which may negatively impact performance in English Zhao et al. (2024); Aggarwal et al. (2024). Few techniques have been proposed to bridge this gap, such as incorporating better pre-training data in various languages and improving base tokenizers Xu et al. (2024); Dagan et al. (2024). However, most of these changes need to be implemented in the pre-training stage, which demands extensive data and computational resources, making it practically unfeasible in many scenarios Brown et al. (2020). Consequently, the most well-studied technique involves fine-tuning models for specific languages and tasks. Instruction fine-tuning (IFT) has become a popular technique to enhance the performance of language models in specific languages. This method combines the benefits of both the pre-training, fine-tuning, and prompting paradigms Wei et al. (2021).
@@
Sample diversity is essential for effective instruction tuning in multilingual datasets. Many recent datasets have been generated by translating English content into other languages or by employing self-instruct techniques based on seed prompts Li et al. (2023); Taori et al. (2023). However, both methods can limit diversity. Machine translation may result in the loss of semantic nuance (Baroni and Bernardini, 2006), while self-instruct approaches often yield repetitive and homogeneous samples Wang et al. (2022). This highlights the critical need for datasets that encompass a wide range of diverse samples.
@@
In this paper, we present a novel recipe for creating a multilingual synthetic instruction tuning dataset, sPhinX.  It comprises 1.8M instruction-response pairs in 51 languages, derived by augmenting the Orca instruction tuning dataset samplesMukherjee et al. (2023) through Selective Translated Augmentation using GPT-4 Achiam et al. (2023). We assess the effectiveness of sPhinX by fine-tuning two models — Phi-3-small and Mistral-7B — across a range of evaluation benchmarks that test various language model capabilities across discriminative and generative tasks. We compare models fine-tuned on sPhinX with those using other synthetic multilingual instruction tuning datasets like Aya Üstün et al. (2024), Multilingual Alpaca Taori et al. (2023), and Bactrian Li et al. (2023) and observe significant performance gains across languages. We also compare our proposed translation strategy with translating the entire instruction using Azure Translator API, as is done with the popular multilingual synthetic IFT datasets to demonstrate the efficacy of our approach.
@@
  • We introduce a novel approach to generate synthetic data for multilingual instruction tuning by Selective Translated Augmentation of the Orca dataset with the assistance of GPT-4 (§3.1)
  • We devise LAnguage-Specific N-shot Guided Instruction Tuning (LANGIT) strategy for enhancing the multilingual capabilities of LLMs (§4)
  • We also conduct extensive instruction tuning experiments on various multilingual instruction tuning datasets to evaluate generalizability in multilingual settings (§6).
  • We plan to release a subset of the augmented dataset by applying our strategy to the OpenOrca  [^1] https://huggingface.co/datasets/Open-Orca/OpenOrca dataset (Lian et al., 2023) (Open-sPhinX) as well.
@@ 2.1 Multilingual Instruction fine-tuning
Early studies focused on fine-tuning pre-trained models on a variety of languages through data augmentation for a single task Hu et al. (2020); Longpre et al. (2021); Asai et al. (2022). Currently, the approach has shifted to fine-tuning these models using a wide variety of tasks Longpre et al. (2023); Ouyang et al. (2022). Models such as BLOOMZ Muennighoff et al. (2022) and mT0 Muennighoff et al. (2022) make significant strides in improving the multilingual performance of decoder-based models Ahuja et al. (2023). There have been multiple multilingual instruction datasets and models proposed such as Bactrian (Li et al., 2023), AYA (Üstün et al., 2024), polyLM (Wei et al., 2023b) after BLOOMZ and mT0 Muennighoff et al. (2023). However, these models still do not perform as well as English in other languages, with the gap being huge for low-resource languages and languages written in scripts other than the Latin script  Ruder et al. (2021); Ahuja et al. (2023); Asai et al. (2024); Ahuja et al. (2024). In this work, we aim to narrow the performance gap by introducing a strategy for creating datasets for multilingual instruction tuning and recipes for fine-tuning, which we will discuss in the following sections.
Figure 1: The figure above illustrates pipelines for sPhinX data creation using Selective Translated Augmentation and Multilingual Instruction Tuning using LANGIT strategy.
Figure 2: The figure compares the Translation API with Selective Translated Augmentation. The Translation API translates the entire input into Russian, while the Selective strategy localizes only necessary components. Here, system and user prompts are translated, but the input question and assistant’s response remain in the original language, preserving structure and intent.
@@ 2.2 Multilingual Synthetic Data Generation
Most instruction-tuning datasets across multiple languages typically focus on general tasks rather than specific reasoning capabilities. Although datasets like Orca Mukherjee et al. (2023) and Orca 2 Mitra et al. (2023) exist in English, they highlight a prevalent issue: current methods often prioritize style imitation over leveraging the reasoning abilities found in large foundation models (LFMs). The Orca dataset addresses this by imitating rich signals from GPT-4, including explanation traces and step-by-step thought processes Wei et al. (2023a), guided by assistance from ChatGPT. In order to create multilingual datasets, researchers commonly use translation APIs or LLMs to translate English-specific datasets into target languages. For example, the Bactrian dataset Li et al. (2023) translates Alpaca and Dolly instructions into 52 languages using the Google Translator API and generates outputs with GPT-3.5 turbo. Another example is of this work  Lai et al. (2024), which also utilizes Google Translation API for translating their source datasets. Our dataset approach aims to tackle these challenges by selectively translating only the essential portions of multilingual inputs. This strategy not only preserves semantic information but also accommodates diverse linguistic contexts, thereby enhancing the overall quality and applicability of instruction-tuning datasets across languages. In our work, we also show the pitfalls of training with data generated only using the Translation APIs such Google Translate or Azure Translate.
@@
In this section, we describe our dataset construction methodology (§3.1), dataset filtering, and cleaning pipelines (§3.2).
@@ 3.1 Dataset Construction
Inspired by (Mukherjee et al., 2023)’s work, we utilized the 1M GPT-4 generated instruction-response pairs from Orca and constructed our own dataset along similar lines using Selective Translated Augmentation into 50 different languages with the help of GPT-4222GPT-4 inference hyper-parameters in Azure OpenAI interface set as: temperature=0.0. We categorize them into three groups: high-resource, mid-resource, and low-resource languages as outlined in Table 7. For high-resource languages, we randomly sample 100k instruction-response pairs from the Orca 1M dataset and generate the responses from GPT-4 with Selective Translated Augmentation as shown in Figure 2. Similarly, we leverage the same strategy for medium and low-resource languages by sampling 50k and 25k pairs respectively. Although GPT-4 performs competitively with commercial translation systems (Google Translate & Bing Translate) it still lags on medium and low resource languages Jiao et al. (2023); Hendy et al. (2023). Furthermore, as highlighted in (Chang et al., 2023; Lin et al., 2023; Xia et al., 2024), fine-tuning with a large set of samples from medium and low-resource languages might lead to catastrophic forgetting of high-resource languages. Therefore, we deliberately create fewer samples for medium and low-resource languages than for high-resource ones. Besides, Shaham et al. (2024) also demonstrates that a small number of multilingual training samples is sufficient to significantly boost multilingual performance, validating our approach of using fewer samples from medium- and low-resource languages.
@@
A fundamental problem with using an off-the-shelf translation API is the lack of semantic and task awareness, in addition to translationese Baroni and Bernardini (2006), which can result in poor quality training data. Consider for example the task of Machine Translation as part of the instruction, wherein the language of the source sentence should be retained. However, an off-the-shelf API, without task awareness, would translate it, resulting in an ambiguous instruction. To mitigate this issue, we used GPT-4 to augment the instructions using Selective Translated Augmentation, so that task-specific components of instruction responses are translated into the appropriate language without changing the semantic meaning. Figure 12 illustrates this with concrete examples. The first example demonstrates the aforementioned translation inconsistency issue for an instruction asking for a French equivalent of an English phrase. The second example demonstrates the consequence of direct translations in the m-Alpaca dataset: wherein the translation of the task input results in the task being ill-defined based on the instructions. As demonstrated, our proposed Selective Translated Augmentation method is able to keep the semantic information of the task intact while translating the instructions. For the exact prompt, please refer to Figure 4 in the Appendix.
3.2 Dataset Filtering and Quality Assessment
@@
After creating the dataset, we filtered out samples where GPT-4 failed to generate a response. The final dataset comprised 1.8 million samples in 51 languages(Table: 15), divided into three subsets: Train, Validation, and Few-shot. Each language’s dataset was partitioned to ensure that the Validation and Few-shot sets contained 2,000 and 1,000 samples, respectively, while the Train set included the remaining data. This approach guarantees consistent distributions across languages in the Validation and Few-shot sets, ensuring equitable representation regardless of the training distribution. The final split ratio for Train, Validation, and Few-shot sets was 92:5.3:2.7.
We also conducted a small-scale quality assessment of the generated data for languages such as Bengali, Hindi, German, Turkish, and Tamil. The researchers and engineers in our organization, who are native speakers of these languages, evaluated the data on the basis of coherence, fluency, and information retention. Our findings indicate that the generated dataset is moderate to high quality.
3.3 Sample Diversity in sPhinX
Unlike prior multilingual datasets such as Bactrian and m-Alpaca, which translate a fixed set of instruction-response pairs into multiple languages, sPhinX ensures diversity by sampling unique subsets of instruction-response pairs for each language.
For instance, Bactrian is constructed from 67k English instruction-response pairs (Alpaca + Dolly) and translated into 52 languages, resulting in identical samples across all languages. In contrast, sPhinX samples from 1M GPT-4-generated instruction-response pairs, ensuring that no two languages share the exact same subset.
Mathematically, the probability that all samples in one language dataset $A$ are also in another language dataset $B$, when sampled without replacement from a larger dataset $D$, is given by:
$P(\text{all A in B})\approx\left(\frac{m}{N}\right)^{n}=\left(\frac{100{,}000}{1{,}000{,}000}\right)^{20{,}000}$
where:
  • $N=1{,}000{,}000$ (Total samples in sPhinX),
  • $n=20{,}000$ (Samples in language $A$),
  • $m=100{,}000$ (Samples in language $B$).
Since the exponential term results in an extremely small probability, this confirms that no two languages have identical instruction-response sets in sPhinX.
To further enhance diversity, we apply Selective Translated Augmentation, translating 10% of samples for high-resource languages, 5% for mid-resource languages, and 2.5% for low-resource languages. This ensures that translated content varies across languages, preventing uniformity.
Additionally, code-switching naturally emerges from this augmentation process, further increasing linguistic diversity. Compared to Aya, which exhibits moderate variation across task instructions, sPhinX introduces greater sample diversity by leveraging a larger and more heterogeneous seed set (Mukherjee et al., 2023) and selective augmentation strategy. Exploring code-switching phenomena would be an interesting task in these synthetically generated datasets, but currently that is out-of-scope of for this work.
4 LANGIT
Following Instruction Tuning strategies of  (Longpre et al., 2023) and also taking inspiration from (Min et al., 2022), we devise Language-Specific N-shot Guided Instruction fine-tuning (LANGIT) Figure: 1. This method aims to improve the model’s ability to follow instructions by augmenting training examples with additional context from a set of few-shot examples in the same language. This added context helps guide the model, enabling it to learn more effectively from the provided examples.
For each training example, we begin by sampling a number of few-shot examples, which are instruction-response pairs in the same language. The number of few-shot examples $N$ is determined probabilistically, with a 30% chance of selecting no few-shot examples, a 20% chance of selecting one, and gradually lower probabilities for higher numbers of few-shots. The maximum number of few-shot examples we sample is six, due to constraints imposed by the model’s context length (8192 tokens) and the typically higher tokenization length in languages other than English.
Once the number of few-shot examples is determined, they are prepended to the main training example, forming an augmented input. This augmented input is then fed into the model for instruction tuning. The purpose of this approach is to expose the model to additional examples of different tasks, helping it generalize better to new tasks in the same language.
We performed experiments to analyze how the model performs on each dataset when fine-tuned using the LANGIT strategy detailed in the next section (§6). Additionally, we fine-tuned the models on the sPhinX dataset without using LANGIT to provide a baseline for comparison. To assess the effectiveness of each instruction-tuning dataset on an equal scale, we conducted a comparative analysis of model performance on different benchmarks, fine-tuning each model on approximately 8 billion tokens per dataset using the LANG strategy.
This fine-tuning strategy is consistently applied across datasets for both the Phi-3-small and Mistral-7B base models. A comparison of token lengths across different datasets is provided in Table 6, showing the average token lengths as tokenized by the Phi-3-small model.